### v2

@@ 1 Introduction
Figure 1: Hatano-Nelson model and its features. (a) Schematic illustration of the model showing the localization of bulk modes under OBCs on the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$). (b) Under PBCs, the eigenvalues trace an ellipse in the complex plane (green), and the eigenstates are extended. In contrast, the OBC spectrum (purple) drastically deviates from the PBC spectrum exhibiting purely real eigenvalues and localized eigenmodes manifesting the NH skin effect. (c) The real (solid) and imaginary (dashed) parts of the PBC energies as a function of $k$. The gray and black crosses in (b) and (c) indicate the left and right movers in the dispersion, respectively. $J_{L}=3/2$ and $J_{R}=1/2$ in (b) and (c).
@@ 2 The NH skin effect: The basics
We start this perspective with a paradigmatic one-dimensional (1D) example, as to discuss several general properties we believe are well-known, yet important to understand the NH skin effect in 1D systems.
@@ 2.1 The Hatano-Nelson model
with $J_{L}\,(J_{R})\in\mathbb{R}$ the nearest-neighbor (NN) hopping parameter to the left (right), and $c^{\dagger}_{n}$ ($c_{n}$) creating (annihilating) an excitation on site $n$. This model is NH when $J_{L}\neq J_{R}$. As a 1D single-band model, the energy under PBCs $E(k)$ corresponds to the Bloch Hamiltonian $H(k)$, and reads $E(k)=(J_{L}+J_{R})\cos k+i(J_{L}-J_{R})\sin k$. As seen in Fig. 1(b), $E(k)$ forms an ellipse and winds around the origin in the complex energy plane as a function of $k$ in the (counter)clockwise direction, when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$). As such, it is possible to define a spectral winding number $w$ as [27, 29, 28]
@@
The winding number can only change when $E(k)=0$ for some $k$, i.e., when $|J_{L}|=|J_{R}|$. Turning to OBCs, it is straightforward to see from the Hamiltonian in Eq. (1), that all eigenstates satisfying $H\ket{\Psi_{R}}=E\ket{\Psi_{R}}$ are exponentially localized on the boundary on the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$) as shown in Fig. 1(a). This phenomenon of a piling up of a macroscopic number of states on the boundary is the NH skin effect.
We find there is an intuitive and enlightening way to understand the connection between the sign of the spectral winding number and the boundary to which the skin states localize. When the winding number is negative (positive), the eigenvalues wind in the (counter)clockwise direction. As the PBC spectrum of the Hatano-Nelson model is centered around zero energy, half of the spectrum is in the positive imaginary plane. That means that for a positive winding number, the right (left) movers with group velocity $v_{k}=\mathop{\mathrm{Re}}[\partial_{k}E(k)]>0$ ($<0$) have negative (positive) imaginary energy, as illustrated in Figs. 1(b,c). In the long-time limit, the left movers thus dominate the behavior of the system, and one sees that all NH skin states pile up on the left boundary of the system. The reverse holds for negative winding number. A similar argument is presented in Ref. 37.
@@ 2.2 Localization of skin states
When considering eigenstates of a NH system, one should take care of the appropriate biorthogonalization [38]: A NH system does not only have right eigenstates |$\ket{\Psi_{R}}$, but also distinct left eigenvectors satisfying $\bra{\Psi_{L}}H=E\bra{\Psi_{L}}$ |, equivalent to $H^{\dagger}\ket{\Psi_{L}}=E^{*}\ket{\Psi_{L}}$. By studying the left and right eigenstates separately one can observe the skin effect. Indeed, if |$\ket{\Psi_{R}}$ is localized on one boundary, the corresponding $\bra{\Psi_{L}}$ | is localized to the opposite boundary [39, 40, 35]. As a consequence, the biorthogonal localization of the right and left eigenstate quantified by $\braket{\Psi_{L}}{\Pi_{n}}{\Psi_{R}}$, where $\Pi_{n}=c_{n}^{\dagger}\ket{0}\bra{0}c_{n}$ is the projector onto each site $n$, has weight throughout the system. Thus, skin states have a bulk-state-like fingerprint in the biorthogonal picture, albeit with potentially negative or complex weights.
@@ 3 Non-normality and non-reciprocity as requirement of the NH skin effect
We want to point out that not all NH matrices feature the NH skin effect, and a necessary criterion needs to be satisfied, namely, the NH matrix needs to be non-normal [1], i.e., $[H,H^{\dagger}]\neq 0$. By writing $H=H_{\text{H}}+iH_{\text{A}}$, with $H_{\text{H}}=H_{\text{H}}^{\dagger}$ the Hermitian part and $iH_{\text{A}}=-(iH_{\text{A}})^{\dagger}$ the anti-Hermitian part of $H$, non-normality is equivalent to $[H_{\text{H}},H_{\text{A}}]\neq 0$. If $H_{\text{H}}$ and $H_{\text{A}}$ would commute, $H$ would share an orthogonal eigenbasis with $H_{\text{H}}$ and $H_{\text{A}}$, which prohibits the existence of skin states.
Additionally, a system needs to be non-reciprocal to show the NH skin effect [53]. This can either be checked by analyzing if the systems Green’s function, $\chi(\omega)=-i(\omega\mathbb{I}-H)^{-1}$, satisfies $|\chi(\omega)|\neq|\chi(\omega)|^{T}$, where the modulus of a matrix means the absolute value of each matrix element, or by analyzing the systems’ singular value decomposition (SVD) [53]. In the scattering context discussed in Ref. 53, the requirement of non-reciprocity is equivalent to the requirement that the systems scattering matrix $S$ satisfies $|S|\neq|S|^{T}$. Physically, this means that the response of the system is not invariant under exchanging input and output [54, 55]. While non-normality and non-reciprocity are equivalent for single-band NN models, e.g., the aforementioned Hatano-Nelson model, this is in general not true for single-band systems including longer-range hoppings and multi-band systems [53].
@@
Figure 2: Topological end state and skin state localization. Normalized right eigenstate of (a) topological end state in pink and (b) skin states in blues for the model in Eq. (4). The end state is localized on a single sublattice whereas the skin modes are distributed across both sublattices. (c) The biorthogonal product of the right and left states shows that the end state remains localized while the skin states are delocalized. We fix $t_{1}=3/2,t_{2}=2,$ $\gamma=0.14$, and $N=41$.
@@ 4 Spectral features, topological protection and the NH bulk-boundary correspondence
If a NH matrix is both non-normal and non-reciprocal, its PBC spectrum displays a so-called point gap [53]. A point gap is defined as there being a base point $E_{B}$, which is not crossed by the complex-energy bands, and crossing this point defines a gap closing transition [29]. Indeed, we see in the example of the Hatano-Nelson model that the PBC spectrum has a point gap, cf. Fig. 1(b), and the gap closes as discussed at $|J_{L}|=|J_{R}|$. It has been shown that for the NH skin effect to occur in 1D systems, the PBC spectrum must feature such point gaps [30, 34, 57, 35]. In extension, this means that it is possible to find a non-zero spectral topological invariant for these systems. For example, the spectral winding number in Eq. (2) for arbitrary periodic 1D systems described by the Bloch Hamiltonian $H(k)$, can be generalized to the following form [29]:
@@
We note that if one is interested in the state localization of a specific skin state, one has to select its energy as reference energy $E_{B}$. While it is quite intuitive that all eigenstates are skewed towards the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$) in the Hatano-Nelson model, the spectral winding number is essential to understand the competition between different hopping strengths and ranges in more general models.
@@
These insights regarding the link between a spectral invariant defined under PBCs, and the NH skin effect then leads to the establishment of a new, truly non-Hermitian bulk-boundary correspondence: Having point gaps in the PBC spectrum of 1D models quantified by a non-trivial spectral invariant implies the piling up of states on the boundary of the system under OBCs [30, 34, 33, 57, 35].
@@ 4.1 Beyond spectral winding numbers: Symmetries
There are models, which feature a NH skin effect, while having a zero spectral winding number. In these cases, it is possible to find alternative spectral invariants. For example, the presence of so-called time-reversal dagger symmetry (TRS[60] with the symmetry operator squaring to minus one results in doubly degenerate loops in the PBC spectrum with opposite winding numbers. As a result, half of the skin states accumulate on one boundary and the other half on the other. The spectral invariant is a $\mathbb{Z}_{2}$ invariant, and this type of skin effect is referred to as the $\mathbb{Z}_{2}$-skin effect [30]. In general, symmetries play an important role in the context of the NH skin effect. Indeed, the presence of certain symmetries prohibits the appearance of the NH skin effect. These symmetries include parity-time symmetry [43] and pseudo-Hermitian symmetry [60] in the unbroken phase for any-dimensional systems as well as parity symmetry and TRS with the symmetry operator squaring to one for 1D systems [60].
@@ 6 Skin effect in higher dimensions
[…]
where $c^{\dagger}_{n,m}$ ($c_{n,m}$) creates (annihilates) an excitation on site $(n,m)$ in the 2D lattice, and $J_{L}$ and $J_{R}$ ($t$) are the (next-)nearest neighbor hoppings. To determine the Bloch Hamiltonian, one Fourier transforms to introduce the two Bloch momenta $k_{x}$ and $k_{y}$. Likewise to the GBZ theory, one analytically continues the Bloch momenta, i.e., one replaces the plane waves $(e^{ik_{x}},e^{ik_{y}})$ by the complex functions $(\beta_{x},\beta_{y})$ to find the non-Bloch Hamiltonian $H(\beta_{x},\beta_{y})$. Then, the formalism relies on the principle that all algebraic-geometric information has to be encoded in its characteristic equation
@@
Figure 3: Amoeba formulation in 2D. (a) Illustration of the 2D single-band model with asymmetric hoppings, Eq. (6). (b,c) The plots depict 2D amoeba, where the points $(\log|\beta_{x}|,\log|\beta_{y}|)$ satisfy the characteristic equation Eq. (6). A hole appears in the amoeba for energies outside the OBC spectrum in (b), while in (c), no hole is present for energies within the OBC spectrum. We set $J_{R}=4/5,J_{R}=6/5$, and $t=1/2$.
@@ 9 Applications
Beyond observing the existence and effects of skin states, an important development we want to highlight is the harnessing of this effect for real world applications. For example, an experiment using fiber loops proposes a topological funnel for light [9]. In this setup, two anisotropic 1D chains with opposite winding numbers are coupled to each other so that the skin effect in both chains conspires to act as a funnel resulting in an accumulation of all excitations at the interface. Another proposal to employ the NH skin effect is in the context of NH topological sensors [61]. Here it is shown that coupling the ends of the anisotropic SSH model, cf. Eq. (4), results in an exponential splitting of the eigenvalues, such that perturbing this coupling leads to an exponentially strong response. A crucial ingredient in this context is the presence of a single topological end state, whose right and left eigenvectors localize on opposite ends. This mechanism is exploited in the experimental realization of the NH topological ohmmeter in Ref. 89. Other NH sensing protocols exist, such as in Ref. 90.

### v3

@@ 1 Introduction
Figure 1: Hatano-Nelson model and its features. (a) Schematic illustration of the model showing the localization of bulk modes under OBCs on the right (left) when $|J_{L}|<|J_{R}|$ ($|J_{R}|$). (b) Under PBCs, the eigenvalues trace an ellipse in the complex plane (green), and the eigenstates are extended. In contrast, the OBC spectrum (purple) drastically deviates from the PBC spectrum exhibiting purely real eigenvalues and localized eigenmodes manifesting the NH skin effect. (c) The real (solid) and imaginary (dashed) parts of the PBC energies as a function of $k$. The gray and black crosses in (b) and (c) indicate the left and right movers in the dispersion, respectively. $J_{L}=3/2$ and $J_{R}=1/2$ in (b) and (c).
@@ 2 The NH skin effect: The basics
We start this perspective with a paradigmatic one-dimensional (1D) example, to discuss several general properties we believe are well-known, yet important to understand the NH skin effect in 1D systems.
@@ 2.1 The Hatano-Nelson model
with $J_{L}\,(J_{R})\in\mathbb{R}$ the nearest-neighbor (NN) hopping parameter to the left (right), and $c^{\dagger}_{n}$ ($c_{n}$) creating (annihilating) an excitation on site $n$. This model is NH when $J_{L}\neq J_{R}$. As a 1D single-band model, the energy under PBCs $E(k)$ corresponds to the Bloch Hamiltonian $H(k)$, and reads $E(k)=(J_{L}+J_{R})\cos k+i(J_{L}-J_{R})\sin k$. As seen in Fig. 1(b), $E(k)$ forms an ellipse and winds around the origin in the complex energy plane as a function of $k$ in the (counter)clockwise direction, when $|J_{L}|<|J_{R}|$ ($|J_{R}|$). As such, it is possible to define a spectral winding number $w$ as [27, 29, 28]
@@
[…]
We find that there is an intuitive and enlightening way to understand the connection between the sign of the spectral winding number and the boundary to which the skin states localize. When the winding number is negative (positive), the eigenvalues wind in the (counter)clockwise direction. As the PBC spectrum of the Hatano-Nelson model is centered around zero energy, half of the spectrum is in the positive imaginary plane. That means that for a positive winding number, the right (left) movers with group velocity $v_{k}=\mathop{\mathrm{Re}}[\partial_{k}E(k)]>0$ (<0<0) have negative (positive) imaginary energy, as illustrated in Figs. 1(b,c). In the long-time limit, the left movers thus dominate the behavior of the system, and one sees that all NH skin states pile up on the left boundary of the system. The reverse holds for negative winding number. A similar argument is presented in Ref. 37.
@@ 2.2 Localization of skin states
[…]
@@ 3 Non-normality and non-reciprocity as requirement of the NH skin effect
We want to point out that not all NH matrices feature the NH skin effect, and a necessary criterion needs to be satisfied, namely, the NH matrix needs to be non-normal [1], i.e., $H,H^{\dagger}]\neq 0$. By writing $H=H_{\text{H}}+iH_{\text{A}}$, with $H_{\text{H}}=H_{\text{H}}^{\dagger}$ the Hermitian part and $iH_{\text{A}}=-(iH_{\text{A}})^{\dagger}$ the anti-Hermitian part of $H$, non-normality is equivalent to $H_{\text{H}},H_{\text{A}}]\neq 0$. If $H_{\text{H}}$ and $H_{\text{A}}$ would commute, $H$ would share an orthogonal eigenbasis with $H_{\text{H}}$ and $H_{\text{A}}$, which prohibits the existence of skin states.
Additionally, a system needs to be non-reciprocal to show the NH skin effect [53]. This can either be checked by analyzing if the systems Green’s function, $\chi(E)=-i(E\,\mathbb{I}-H)^{-1}$, satisfies $|\chi(E)|\neq|\chi(E)|^{T}$, where the modulus of a matrix means the absolute value of each matrix element, or by analyzing the systems’ singular value decomposition (SVD) [53]. In the scattering context discussed in Ref. 53, the requirement of non-reciprocity is equivalent to the requirement that the systems scattering matrix $S$ satisfies $|S|\neq|S|^{T}$. Physically, this means that the response of the system is not invariant under exchanging input and output [54, 55]. While non-normality and non-reciprocity are equivalent for single-band NN models, e.g., the aforementioned Hatano-Nelson model, this is in general not true for single-band systems including longer-range hoppings and multi-band systems [53].
@@
Figure 2: Topological end state and skin state localization. Normalized right eigenstate of (a) topological end state in pink and (b) skin states in different shades of blue for the model in Eq. (4). The end state is localized on a single sublattice whereas the skin modes are distributed across both sublattices. (c) The biorthogonal product of the right and left states shows that the end state remains localized while the skin states are delocalized. We fix $t_{1}=3/2,t_{2}=2,$ $\gamma=0.14$, and $N=41$.
@@ 4 Spectral features, topological protection and the NH bulk-boundary correspondence
If a NH matrix is both non-normal and non-reciprocal, its PBC spectrum displays a so-called point gap [53]. A point gap is defined as there being a base point $E_{B}$, which is not crossed by the complex-energy bands, and crossing this point defines a gap closing transition [29]. Indeed, we see in the example of the Hatano-Nelson model that the PBC spectrum has a point gap, cf. Fig. 1(b), and the gap closes as discussed at $|J_{L}|=|J_{R}|$. It has been shown that for the NH skin effect to occur in 1D systems, the PBC spectrum must feature such point gaps [30, 34, 57, 35]. By extension, this means that it is possible to find a non-zero spectral topological invariant for these systems. For example, the spectral winding number in Eq. (2) for arbitrary periodic 1D systems described by the Bloch Hamiltonian $H(k)$, can be generalized to the following form [29]:
@@
We note that if one is interested in the state localization of a specific skin state, one has to select its energy as reference energy $E_{B}$. While it is quite intuitive that all eigenstates are skewed towards the right (left) when $|J_{L}|<|J_{R}|$ ($|J_{R}|$) in the Hatano-Nelson model, the spectral winding number is essential to understand the competition between different hopping strengths and ranges in more general models.
@@
These insights regarding the link between a spectral invariant defined under PBCs, and the NH skin effect then lead to the establishment of a new, truly non-Hermitian bulk-boundary correspondence: Having point gaps in the PBC spectrum of 1D models quantified by a non-trivial spectral invariant implies the piling up of states on the boundary of the system under OBCs [30, 34, 33, 57, 35].
@@ 4.1 Beyond spectral winding numbers: Symmetries
There are models, which feature a NH skin effect, while having a zero spectral winding number. In these cases, it is possible to find alternative spectral invariants. For example, the presence of the so-called time-reversal dagger symmetry (TRS[60] with the symmetry operator squaring to minus one results in doubly degenerate loops in the PBC spectrum with opposite winding numbers. As a result, half of the skin states accumulate on one boundary and the other half on the other. The spectral invariant is a $\mathbb{Z}_{2}$ invariant, and this type of skin effect is referred to as the $\mathbb{Z}_{2}$-skin effect [30]. In general, symmetries play an important role in the context of the NH skin effect. Indeed, the presence of certain symmetries prohibits the appearance of the NH skin effect. These symmetries include parity-time symmetry [43] and pseudo-Hermitian symmetry [60] in the unbroken phase for any-dimensional systems as well as parity symmetry and TRS with the symmetry operator squaring to one for 1D systems [60].
@@ 6 Skin effect in higher dimensions
$\quad\ +t\left.\left(c^{\dagger}_{n,m}c_{n+1,m+1}+c^{\dagger}_{n,m+1}c_{n+1,m}+\text{H.c.}\right)\right],$ (5)
where $c^{\dagger}_{n,m}$ ($c_{n,m}$) creates (annihilates) an excitation on site $(n,m)$ in the 2D lattice, and $J_{L}$ and $J_{R}$ ($t$) are the (next-)nearest neighbor hoppings. To determine the Bloch Hamiltonian, one Fourier transforms to introduce the two Bloch momenta $k_{x}$ and $k_{y}$. As in the GBZ theory, one analytically continues the Bloch momenta, i.e., one replaces the plane waves $(e^{ik_{x}},e^{ik_{y}})$ by the complex functions $(\beta_{x},\beta_{y})$ to find the non-Bloch Hamiltonian $H(\beta_{x},\beta_{y})$. Then, the formalism relies on the principle that all algebraic-geometric information has to be encoded in its characteristic equation
@@
Figure 3: Amoeba formulation in 2D. (a) Illustration of the 2D single-band model with asymmetric hoppings, Eq. (6). (b,c) The plots depict 2D amoeba, where the points $(\log|\beta_{x}|,\log|\beta_{y}|)$ satisfy the characteristic equation Eq. (6). A hole appears in the amoeba for energies outside the OBC spectrum in (b), while in (c), no hole is present for energies within the OBC spectrum. We set $J_{R}=4/5,J_{L}=6/5$, and $t=1/2$.
@@ 9 Applications
Beyond observing the existence and effects of skin states, an important development we want to highlight is the harnessing of this effect for real world applications. For example, an experiment using fiber loops proposes a topological funnel for light [9]. In this setup, two anisotropic 1D chains with opposite winding numbers are coupled to each other so that the skin effect in both chains acts as a funnel resulting in an accumulation of all excitations at the interface. Another proposal to employ the NH skin effect is in the context of NH topological sensors [61]. Here it is shown that coupling the ends of the anisotropic SSH model, cf. Eq. (4), results in an exponential splitting of the eigenvalues, such that perturbing this coupling leads to an exponentially strong response. A crucial ingredient in this context is the presence of a single topological end state, whose right and left eigenvectors localize on opposite ends. This mechanism is exploited in the experimental realization of the NH topological ohmmeter in Ref. 89. Other NH sensing protocols exist, such as in Ref. 90.